OpenStudy (madgirlwithabluebox):

Will fan and medal: HELP!! What is an equation of a parabola with the given vertex and focus? vertex: (–2, 5); focus: (–2, 6)

OpenStudy (campbell_st):

start by drawing a diagram  so the equation is in the form \[(y - k)^2 = 4a(x - h)\] (h, k) is the vertex a is the focal length ... of the distance between the vetrex and focus

OpenStudy (madgirlwithabluebox):

so \[(y-5)^2=4(1)(x+2) \] ?

OpenStudy (rizags):

no, campbell's diagram and equation are in the incorrect form.

OpenStudy (madgirlwithabluebox):

well i already submitted it so its cool..could you help me work out the correct answer anyway so i know how to find it next time?

OpenStudy (rizags):

ok, so the true equation of a parabola in the form you described, is \[4a(y-k)=(x-h)^2\]

OpenStudy (rizags):

yea just plug in everything and thats the equation

OpenStudy (rizags):

got it? Medal please